stormyfurr:

2/3=5/x A. 2/15 B. 6 C. 15/2 D. 8

stormyfurr:

Please help me.

mxddi3:

do yk how to cross multiply?

stormyfurr:

A little

stormyfurr:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @stormyfurr A little \(\color{#0cbb34}{\text{End of Quote}}\) Well, not that much.

mxddi3:

ok so to get x, you can cross multiply \[\frac{ 2 }{ 3 }= \frac{ 5 }{x }\] first step is to do 5x3=? and then 2(x)=?

mxddi3:

solve those 2 for me please

stormyfurr:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @mxddi3 solve those 2 for me please \(\color{#0cbb34}{\text{End of Quote}}\) Will do

mxddi3:

once you get those, tell me what yu got

stormyfurr:

I say that the answer is \[\frac{ 15 }{ 2 }\]

mxddi3:

mhmm, good job

stormyfurr:

Thanks

mxddi3:

np